### v1

@@ 4 A preconditioned correction equation, MINRES, and practical issues
Recall (3.16)–(3.18), and write $\widetilde{W}_{1}=W_{1}$ in (3.1) for $\tilde{u}_{1}=u_{1}$ and $\tilde{v}_{1}=v_{1}$. Then the normalized eigenvectors of $\widetilde{B}_{1}=B_{1}=W_{1}B_{1}^{\prime}W_{1}^{T}$ defined in (3.3) associated with the eigenvalues ±$\pm\sigma_{2}-\tau,\dots,\pm\sigma_{N}-\tau$ and $-\tau$ are $\frac{1}{\sqrt{2}}[u_{i}^{T},\pm v_{i}^{T}]^{T},\ i=2,\ldots,N$ and $[u^{T}_{i},\mathbf{0}]^{T}$, $i=N+1,N+2,\ldots,M$, respectively; see (1.1). Therefore, those smallest eigenvalues in magnitude of $B_{1}$ correspond to the singular values of $A$ closest to $\tau$. Notice that for $\tilde{u}_{1}\not=u_{1}$ and $\tilde{v}_{1}\not=v_{1}$, since $\widetilde{B}_{1}\rightarrow B_{1}$ as $\tilde{u}_{1}\rightarrow u_{1}$ and $\tilde{v}_{1}\rightarrow v_{1}$, the eigenvalues and eigenvectors (rigorously speaking, the eigenspaces associated with each of the well isolated clusters of eigenvalues) of $\widetilde{B}_{1}$ converge to those of $B_{1}$ correspondingly. As a result, for $\tilde{u}_{1}\rightarrow u_{1}$ and $\tilde{v}_{1}\rightarrow v_{1}$, if $B_{1}$ has clustered smallest eigenvalues in magnitude well separated from the others, then $\widetilde{B}_{1}$ has the same number of clustered smallest eigenvalues in magnitude well isolated with the others.
@@
The preceding convergence results show that MINRES may converge very slowly when $\sigma_{1}$ is clustered with some other singular values of $A$, which corresponds to the case that the clustered smallest eigenvalues in magnitude of $\widetilde{B}_{1}$ are well separated from the others. In this case, it is vital to precondition the correction equation (2.7) in order to accelerate MINRES. Theoretically, this can be realized by annihilating the effects of those smallest eigenvalues in magnitude of $\widetilde{B}_{1}$. Strikingly, as JDSVD proceeds, the searching subspaces $\mathcal{U}$ and $\mathcal{V}$ generally contain increasingly more accurate approximations to the singular vectors of $A$ associated with several singular values closest to $\tau$. Based on the last paragraph, this remarkable property indicates that JDSVD provides rich information on the eigenvectors of $B_{1}$ and thus those of $\widetilde{B}_{1}$ associated with those smallest eigenvalues in magnitude. It, therefore, makes us possible to fully exploit such information to precondition the correction equation effectively so as to solve it by MINRES more efficiently, which leads to our inner preconditioned JDSVD (IPJDSVD) method. We will show that such preconditioning is equivalent to forming a preconditioned correction equation, that makes full use of all the approximate left and right singular vectors associated with the clustered singular values nearest to $\tau$. If the clustered singular values closest to $\tau$ are well isolated with the other singular values of $A$, those smallest eigenvalues in magnitude of the preconditioned coefficient matrix are well separated from the other eigenvalues, so that the convergence of MINRES is improved substantially.
@@
Section 4.1 is devoted to a detailed derivation of the preconditioned correction equation; section 4.2 presents convergence results on MINRES for solving it; section 4.3 settles down the practical issue of how to adaptively form the preconditioned correction equation at each outer iteration.
@@
(4.1) $|\sigma_{1}-\tau|\lesssim\dots\lesssim|\sigma_{m}-\tau|\ll|\sigma_{m+1}-\tau|\leq\dots\leq|\sigma_{N}-\tau|.$
From (1.1), we denote by
[…]
@@
At the $k$th outer iteration of JDSVD, suppose that $k>m$. Recall from section 2 that $(\theta_{i},c_{i},d_{i})$, $i=1,\dots,k$ are the singular triplets of $H$ in (2.2), labeled as $|\theta_{1}-\tau|\leq|\theta_{2}-\tau|\leq\dots\leq|\theta_{k}-\tau|$. We compute the $k$ Ritz approximations
(4.3) $(\theta_{i},\tilde{u}_{i},\tilde{v}_{i})=(\theta_{i},\widetilde{U}c_{i},\widetilde{V}d_{i}),\qquad i=1,\dots,k,$
@@
[…]
@@ 4.1 A preconditioned correction equation
By (2.1) and (2.4), from (4.4) we have
[…]
(4.6) $\begin{bmatrix}I-\widetilde{U}_{m}\widetilde{U}_{m}^{T}&\\ &I-\widetilde{V}_{m}\widetilde{V}_{m}^{T}\end{bmatrix}\begin{bmatrix}-\tau I&A\\ A^{T}&-\tau I\end{bmatrix}\begin{bmatrix}s\\ t\end{bmatrix}=-r,$
@@
and substituting this relation into (4.6), we obtain
(4.7) $\begin{bmatrix}I-\widetilde{U}_{m}\widetilde{U}_{m}^{T}\!\!\!\!&\\ &\!\!\!\!I-\widetilde{V}_{m}\widetilde{V}_{m}^{T}\end{bmatrix}\begin{bmatrix}-\tau I&A\\ A^{T}&-\tau I\end{bmatrix}\begin{bmatrix}I-\widetilde{U}_{m}\widetilde{U}_{m}^{T}\!\!\!\!&\\ &\!\!\!\!I-\widetilde{V}_{m}\widetilde{V}_{m}^{T}\end{bmatrix}\begin{bmatrix}s\\ t\end{bmatrix}=-r-r_{\mathrm{tail}}$
with
(4.8) $r_{\mathrm{tail}}$ $=$ $\begin{bmatrix}I-\widetilde{U}_{m}\widetilde{U}_{m}^{T}&\\ &I-\widetilde{V}_{m}\widetilde{V}_{m}^{T}\end{bmatrix}\begin{bmatrix}-\tau I&A\\ A^{T}&-\tau I\end{bmatrix}\begin{bmatrix}\widetilde{U}_{m}\widetilde{U}_{m}^{T}&\\ &\widetilde{V}_{m}\widetilde{V}_{m}^{T}\end{bmatrix}\begin{bmatrix}s\\ t\end{bmatrix}$
@@
$=$ $\begin{bmatrix}&A\widetilde{V}_{m}-\widetilde{U}_{m}\Theta_{m}\\ A^{T}\widetilde{U}_{m}-\widetilde{V}_{m}\Theta_{m}\end{bmatrix}\begin{bmatrix}\widetilde{U}_{m}^{T}s\\ \widetilde{V}_{m}^{T}t\end{bmatrix},$
where the last equality comes from $\Theta_{m}=\widetilde{U}_{m}^{T}A\widetilde{V}_{m}$.
@@
From (2.7), the first elements of $\widetilde{U}_{m}^{T}s$ and $\widetilde{V}_{m}^{T}t$ are $\tilde{u}_{1}^{T}s=0$ and $\tilde{v}_{1}^{T}t=0$. Denote $\widetilde{U}_{m}^{\prime}=[\tilde{u}_{2},\dots,\tilde{u}_{m}]$, $\widetilde{V}_{m}^{\prime}=[\tilde{v}_{2},\dots,\tilde{v}_{m}]$ and $\Theta_{m}^{\prime}=[\theta_{2},\dots,\theta_{m}]$. Then (4.8) becomes
(4.9) $r_{\mathrm{tail}}=\begin{bmatrix}&\!\!\!\!A\widetilde{V}_{m}^{\prime}-\widetilde{U}_{m}^{\prime}\Theta_{m}^{\prime}\\ A^{T}\widetilde{U}_{m}^{\prime}-\widetilde{V}_{m}^{\prime}\Theta_{m}^{\prime}&\end{bmatrix}\begin{bmatrix}(\widetilde{U}_{m}^{\prime})^{T}s\\ (\widetilde{V}_{m}^{\prime})^{T}t\end{bmatrix}=\begin{bmatrix}&\!\!\!\!R_{1}\\ R_{2}&\end{bmatrix}\begin{bmatrix}(\widetilde{U}_{m}^{\prime})^{T}s\\ (\widetilde{V}_{m}^{\prime})^{T}t\end{bmatrix},$
@@
where $R_{1}=A\widetilde{V}_{m}^{\prime}-\widetilde{U}_{m}^{\prime}\Theta_{m}^{\prime}$ and $R_{2}=A^{T}\widetilde{U}_{m}^{\prime}-\widetilde{V}_{m}^{\prime}\Theta_{m}^{\prime}$, whose $(i-1)$-th columns are respectively the upper and lower parts of the residual $r^{(i)}=r(\theta_{i},\tilde{u}_{i},\tilde{v}_{i})$ of $(\theta_{i},\tilde{u}_{i},\tilde{v}_{i})$ defined by (2.4), $i=2,\dots,m$. Keep in mind the three properties: (i) the sizes of $\|s\|$ and $\|t\|$ are of the same orders as the errors of $\tilde{u}_{i}$ and $\tilde{v}_{i}$, respectively; (ii) $\|r^{(i)}\|$ is of the same order as the errors of $\tilde{u}_{i},\tilde{v}_{i}$, $i=1,\dots,m$; (iii) $(\widetilde{U}_{m}^{\prime})^{T}s$ and $(\widetilde{V}_{m}^{\prime})^{T}t$ are no more than $s$ and $t$ in size. Therefore, whenever $\|r^{(i)}\|,i=2,\ldots,m$ are comparable to $\|r\|$, that is, $(\theta_{i},\tilde{u}_{i},\tilde{v}_{i}),i=2,\dots,m$ have comparable accuracy to $(\theta_{1},\tilde{u}_{1},\tilde{v}_{1})$, one has $\|r_{\rm tail}\|=\mathcal{O}(\|r\|^{2})$ and thus $r+r_{\rm tail}\approx r$ in (4.7) once $\|r\|$ becomes fairly small.
@@
Omitting $r_{\mathrm{tail}}$ in (4.7), we have now derived a preconditioned correction equation
(4.10) $\begin{bmatrix}I-\widetilde{U}_{m}\widetilde{U}_{m}^{T}&\\ &I-\widetilde{V}_{m}\widetilde{V}_{m}^{T}\end{bmatrix}\begin{bmatrix}-\tau I&A\\ A^{T}&-\tau I\end{bmatrix}\begin{bmatrix}I-\widetilde{U}_{m}\widetilde{U}_{m}^{T}&\\ &I-\widetilde{V}_{m}\widetilde{V}_{m}^{T}\end{bmatrix}\begin{bmatrix}s\\ t\end{bmatrix}=-r,$
@@
whose solution $(s,t)\perp\perp(\widetilde{U}_{m},\widetilde{V}_{m})$ has the same quality as the solution $(s,t)$ to (2.7) provided that $(\theta_{i},\tilde{u}_{i},\tilde{v}_{i})$, $i=2,\dots,m$ have comparable accuracy to $(\theta_{1},\tilde{u}_{1},\tilde{v}_{1})$.
@@
With an approximate solution $(\tilde{s}^{T},\tilde{t}^{T})^{T}$ of (4.10) available, we orthonormalize $\tilde{s}$ and $\tilde{t}$ against $\widetilde{U}$ and $\widetilde{V}$ to obtain the expansion vectors $u_{+}$ and $v_{+}$, and expand the subspaces in a regular way until $(\theta_{1},\tilde{u}_{1},\tilde{v}_{1})$ converges. This forms our basic IPJDSVD. Importantly, the above derivation and analysis indicate that the solution to the preconditioned correction equation expands the searching subspaces as effectively as the solution to the standard correction equation (2.7) does. In other words, IPJDSVD does not spoil the subsequent outer iterations of JDSVD.
4.2 Convergence analysis on MINRES for solving (4.10)
@@
(4.11) $\widetilde{W}_{m}=\begin{bmatrix}\widetilde{P}_{m}&\\ &\widetilde{Q}_{m}\end{bmatrix}\in\mathbb{R}^{(M+N)\times(M+N-2m)}.$
Then the coefficient matrix in (4.10) can be written as
(4.12) $\widetilde{B}_{m}=\widetilde{W}_{m}\widetilde{B}_{m}^{\prime}\widetilde{W}_{m}^{T}\quad\mbox{with}\quad\widetilde{B}_{m}^{\prime}=\begin{bmatrix}-\tau I_{M-m}&\widetilde{P}_{m}^{T}A\widetilde{Q}_{m}\\ \widetilde{Q}_{m}^{T}A^{T}\widetilde{P}_{m}&-\tau I_{N-m}\end{bmatrix}.$
Denote by $\tilde{w}_{j}$ the $j$th approximate solution computed by MINRES for (4.10) with $\tilde{w}_{0}=\bm{0}$, $j=1,2,\ldots,M+N-2m$. Then the corresponding inner residuals are
(4.13) $\qquad r_{\mathrm{in},j}=\widetilde{B}_{m}\tilde{w}_{j}+r,\qquad\qquad j=0,1,\ldots,M+N-2m.$
@@
With $\widetilde{B}_{m}^{\prime}$ defined in (4.12), the solution to the preconditioned correction equation (4.10) is $w=\widetilde{W}_{m}z$ where $z=\widetilde{W}_{m}^{T}w\in\mathbb{R}^{M+N-2m}$ solves
(4.14) $\widetilde{B}_{m}^{\prime}z=-r^{\prime}\qquad\mbox{with}\qquad r^{\prime}=\widetilde{W}_{m}^{T}r.$
If MINRES is applied to (4.14) with the initial guess $\tilde{z}_{0}=\bm{0}$, then the $j$th approximate solution $\tilde{z}_{j}=\widetilde{W}_{m}^{T}\tilde{w}_{j}$ with $\tilde{w}_{j}=\widetilde{W}_{m}\tilde{z}_{j}$ and
(4.15) $\|r_{\mathrm{in},j}\|=\|r^{\prime}_{\mathrm{in},j}\|=\|\widetilde{B}_{m}^{\prime}\tilde{z}_{j}+r^{\prime}\|,\ j=1,2,\ldots,M+N-2m.$
Theorem 4.1 establishes the equivalence of MINRES for solving (4.10) and (4.14). Notice that $r=\mathbf{0}$ when $\widetilde{U}_{m}=U_{m}$ and $\widetilde{V}_{m}=V_{m}$. In this case, we can extend Theorem 3.3 to MINRES for solving (4.10) by supposing any virtual right hand side $r\not=\mathbf{0}$ that satisfies (4.5).
@@
Suppose that $\mathcal{R}(\widetilde{U}_{m})=\mathcal{R}(U_{m})$ and $\mathcal{R}(\widetilde{V}_{m})=\mathcal{R}(V_{m})$ in (4.10) with $U_{m}$ and $V_{m}$ in (4.2) and $\mathcal{R}(\cdot)$ the column space of a matrix. Then (i) if $\tau>(\sigma_{\max}+\sigma_{\max,m+1})/2$ with $\sigma_{\max,m+1}$ the $(m+1)$th largest singular value of $A$, then
(4.16) $\frac{\|r_{\mathrm{in},j}\|}{\|r\|}\leq 2\left(1-\frac{2}{1+\sqrt{(\tau+\sigma_{\max,m+1})/(\tau-\sigma_{\max,m+1})}}\right)^{j};$
(ii) if $\tau<(\sigma_{\min}+\sigma_{\min,m+1})/2$ with $\sigma_{\min,m+1}$ the $(m+1)$th smallest singular value of $A$ and writing $j_{o}=\min\{1,M-N\}$, then
(4.17) $\frac{\|r_{\mathrm{in},j}\|}{\|r\|}\leq 2\left(\frac{\sigma_{\max}}{\tau}\right)^{j_{o}}\left(1-\frac{2}{1+\sqrt{(\sigma_{\max}^{2}-\tau^{2})/(\sigma_{\min,m+1}^{2}-\tau^{2})}}\right)^{\left[\frac{j-j_{o}}{2}\right]};$

### v2

@@ 4 A preconditioned correction equation, MINRES, and practical issues
Recall (25)–(27), and write $\widetilde{W}_{1}=W_{1}$ in (10) for $\tilde{u}_{1}=u_{1}$ and $\tilde{v}_{1}=v_{1}$. Then the normalized eigenvectors of $\widetilde{B}_{1}=B_{1}=W_{1}B_{1}^{\prime}W_{1}^{T}$, defined in (12), associated with the eigenvalues ±$\pm\sigma_{2}-\tau,\dots,\pm\sigma_{N}-\tau$ and $-\tau$ are $\frac{1}{\sqrt{2}}[u_{i}^{T},\pm v_{i}^{T}]^{T},\ i=2,\ldots,N$ and $[u^{T}_{i},\mathbf{0}]^{T}$, $i=N+1,N+2,\ldots,M$, respectively; see (1). Therefore, those smallest eigenvalues in magnitude of $B_{1}$ correspond to the singular values of $A$ closest to $\tau$. Notice that for $\tilde{u}_{1}\not=u_{1}$ and $\tilde{v}_{1}\not=v_{1}$, since $\widetilde{B}_{1}\rightarrow B_{1}$ as $\tilde{u}_{1}\rightarrow u_{1}$ and $\tilde{v}_{1}\rightarrow v_{1}$, the eigenvalues and eigenvectors (rigorously speaking, the eigenspaces associated with each of the well isolated clusters of eigenvalues) of $\widetilde{B}_{1}$ converge to those of $B_{1}$ correspondingly. As a result, for $\tilde{u}_{1}\rightarrow u_{1}$ and $\tilde{v}_{1}\rightarrow v_{1}$, if the clustered smallest eigenvalues of $B_{1}$ are well separated from the others, then $\widetilde{B}_{1}$ has the same number of clustered smallest eigenvalues in magnitude.
@@
The preceding convergence results show that MINRES may converge very slowly when $\sigma_{1}$ is clustered with some other singular values of $A$, which corresponds to the case that the clustered smallest eigenvalues in magnitude of $\widetilde{B}_{1}$ are well separated from the others. In this case, it is vital to precondition the correction equation (9) in order to accelerate MINRES. Theoretically, this can be realized by annihilating the effects of those smallest eigenvalues in magnitude of $\widetilde{B}_{1}$. Strikingly, as JDSVD proceeds, the subspaces $\mathcal{U}$ and $\mathcal{V}$ generally contain increasingly more accurate approximations to the singular vectors of $A$ associated with several singular values closest to $\tau$. Based on the last paragraph, this property indicates that JDSVD provides rich information on the eigenvectors of $B_{1}$ and thus those of $\widetilde{B}_{1}$ associated with the smallest eigenvalues in magnitude. It, therefore, makes us possible to exploit such information to precondition the correction equation effectively so as to solve it by MINRES more efficiently, which leads to our inner preconditioned JDSVD (IPJDSVD) method. We will show that such preconditioning is equivalent to forming a preconditioned correction equation that makes full use of all the approximate left and right singular vectors associated with the clustered singular values nearest to $\tau$. If the clustered singular values closest to $\tau$ are well isolated with the other singular values of $A$, those smallest eigenvalues in magnitude of the preconditioned coefficient matrix are well separated from the other eigenvalues, too, so that the convergence of MINRES is improved substantially.
@@
Subsection 4.1 is devoted to a detailed derivation of the preconditioned correction equation; Section 4.2 presents convergence results on MINRES for solving it; Section 4.3 settles down the practical issue of how to adaptively form the preconditioned correction equation at each outer iteration.
@@
(47) $|\sigma_{1}-\tau|\lesssim\dots\lesssim|\sigma_{m}-\tau|\ll|\sigma_{m+1}-\tau|\leq\dots\leq|\sigma_{N}-\tau|.$
From (1), we denote by
(48) $\Sigma_{m}=\operatorname{diag}\{\sigma_{1},\dots,\sigma_{m}\},\qquad U_{m}=[u_{1},\dots,u_{m}],\qquad V_{m}=[v_{1},\dots,v_{m}].$
@@
At the $k$th outer iteration of JDSVD, suppose that $k>m$. Recall from Section 2 that $(\theta_{i},c_{i},d_{i})$, $i=1,\dots,k$ are the singular triplets of $H$ in (4), labeled as $|\theta_{1}-\tau|\leq|\theta_{2}-\tau|\leq\dots\leq|\theta_{k}-\tau|$. We compute the $k$ Ritz approximations
(49) $(\theta_{i},\tilde{u}_{i},\tilde{v}_{i})=(\theta_{i},\widetilde{U}c_{i},\widetilde{V}d_{i}),\qquad i=1,\dots,k,$
@@
(50) $\Theta_{m}=\operatorname{diag}\{\theta_{1},\dots,\theta_{m}\},\qquad\widetilde{U}_{m}=[\tilde{u}_{1},\dots,\tilde{u}_{m}],\qquad\widetilde{V}_{m}=[\tilde{v}_{1},\dots,\tilde{v}_{m}].$
@@ 4.1 A preconditioned correction equation
By (3) and (6), from (49)–(50) we have
(51) $r\perp\perp\left(\widetilde{U}_{m},\widetilde{V}_{m}\right)\mbox{,\ \ i.e.,\ \ }r=\operatorname{diag}\{I-\widetilde{U}_{m}\widetilde{U}_{m}^{T},I-\widetilde{V}_{m}\widetilde{V}_{m}^{T}\}r.$
Therefore, premultiplying (9) by $\operatorname{diag}\{I-\widetilde{U}_{m}\widetilde{U}_{m}^{T},I-\widetilde{V}_{m}\widetilde{V}_{m}^{T}\}$ yields
(52) $\begin{bmatrix}I-\widetilde{U}_{m}\widetilde{U}_{m}^{T}&\\ &I-\widetilde{V}_{m}\widetilde{V}_{m}^{T}\end{bmatrix}\begin{bmatrix}-\tau I&A\\ A^{T}&-\tau I\end{bmatrix}\begin{bmatrix}s\\ t\end{bmatrix}=-r,$
@@
and substituting this relation into (52), we obtain
(53) $\begin{bmatrix}I-\widetilde{U}_{m}\widetilde{U}_{m}^{T}\!\!\!\!&\\ &\!\!\!\!I-\widetilde{V}_{m}\widetilde{V}_{m}^{T}\end{bmatrix}\begin{bmatrix}-\tau I&A\\ A^{T}&-\tau I\end{bmatrix}\begin{bmatrix}I-\widetilde{U}_{m}\widetilde{U}_{m}^{T}\!\!\!\!&\\ &\!\!\!\!I-\widetilde{V}_{m}\widetilde{V}_{m}^{T}\end{bmatrix}\begin{bmatrix}s\\ t\end{bmatrix}=-r-r_{\mathrm{tail}},$
where, by $\Theta_{m}=\widetilde{U}_{m}^{T}A\widetilde{V}_{m}$,
(54) $r_{\mathrm{tail}}$ $=$ $\begin{bmatrix}I-\widetilde{U}_{m}\widetilde{U}_{m}^{T}&\\ &I-\widetilde{V}_{m}\widetilde{V}_{m}^{T}\end{bmatrix}\begin{bmatrix}-\tau I&A\\ A^{T}&-\tau I\end{bmatrix}\begin{bmatrix}\widetilde{U}_{m}\widetilde{U}_{m}^{T}&\\ &\widetilde{V}_{m}\widetilde{V}_{m}^{T}\end{bmatrix}\begin{bmatrix}s\\ t\end{bmatrix}$
@@
$=$ $\begin{bmatrix}&A\widetilde{V}_{m}-\widetilde{U}_{m}\Theta_{m}\\ A^{T}\widetilde{U}_{m}-\widetilde{V}_{m}\Theta_{m}\end{bmatrix}\begin{bmatrix}\widetilde{U}_{m}^{T}s\\ \widetilde{V}_{m}^{T}t\end{bmatrix}.$
@@
From (9), the first elements of $\widetilde{U}_{m}^{T}s$ and $\widetilde{V}_{m}^{T}t$ are $\tilde{u}_{1}^{T}s=0$ and $\tilde{v}_{1}^{T}t=0$. Denote $\widetilde{U}_{m}^{\prime}=[\tilde{u}_{2},\dots,\tilde{u}_{m}]$, $\widetilde{V}_{m}^{\prime}=[\tilde{v}_{2},\dots,\tilde{v}_{m}]$ and $\Theta_{m}^{\prime}=[\theta_{2},\dots,\theta_{m}]$. Then (54) becomes
(55) $r_{\mathrm{tail}}=\begin{bmatrix}&\!\!\!\!A\widetilde{V}_{m}^{\prime}-\widetilde{U}_{m}^{\prime}\Theta_{m}^{\prime}\\ A^{T}\widetilde{U}_{m}^{\prime}-\widetilde{V}_{m}^{\prime}\Theta_{m}^{\prime}&\end{bmatrix}\begin{bmatrix}(\widetilde{U}_{m}^{\prime})^{T}s\\ (\widetilde{V}_{m}^{\prime})^{T}t\end{bmatrix}=\begin{bmatrix}&\!\!\!\!R_{1}\\ R_{2}&\end{bmatrix}\begin{bmatrix}(\widetilde{U}_{m}^{\prime})^{T}s\\ (\widetilde{V}_{m}^{\prime})^{T}t\end{bmatrix},$
@@
where $R_{1}=A\widetilde{V}_{m}^{\prime}-\widetilde{U}_{m}^{\prime}\Theta_{m}^{\prime}$ and $R_{2}=A^{T}\widetilde{U}_{m}^{\prime}-\widetilde{V}_{m}^{\prime}\Theta_{m}^{\prime}$, whose $(i-1)$-th columns are respectively the upper and lower parts of the residual $r^{(i)}=r(\theta_{i},\tilde{u}_{i},\tilde{v}_{i})$ of $(\theta_{i},\tilde{u}_{i},\tilde{v}_{i})$ defined by (6), $i=2,\dots,m$. Keep in mind the three properties: (i) $\|s\|$ and $\|t\|$ are of the same orders as the errors of $\tilde{u}_{i}$ and $\tilde{v}_{i}$, respectively; (ii) $\|r^{(i)}\|$ is of the same order as the errors of $\tilde{u}_{i},\tilde{v}_{i}$, $i=1,\dots,m$; (iii) $(\widetilde{U}_{m}^{\prime})^{T}s$ and $(\widetilde{V}_{m}^{\prime})^{T}t$ are no more than $s$ and $t$ in size. Therefore, whenever $\|r^{(i)}\|,i=2,\ldots,m$ are comparable to $\|r\|$, one has $\|r_{\rm tail}\|=\mathcal{O}(\|r\|^{2})$ and thus $r+r_{\rm tail}\approx r$ in (53) once $\|r\|$ becomes fairly small.
@@
Omitting $r_{\mathrm{tail}}$ in (53), we have now derived a preconditioned correction equation:
(56) $\begin{bmatrix}I-\widetilde{U}_{m}\widetilde{U}_{m}^{T}&\\ &I-\widetilde{V}_{m}\widetilde{V}_{m}^{T}\end{bmatrix}\begin{bmatrix}-\tau I&A\\ A^{T}&-\tau I\end{bmatrix}\begin{bmatrix}I-\widetilde{U}_{m}\widetilde{U}_{m}^{T}&\\ &I-\widetilde{V}_{m}\widetilde{V}_{m}^{T}\end{bmatrix}\begin{bmatrix}s\\ t\end{bmatrix}=-r,$
@@
whose solution $(s,t)\perp\perp(\widetilde{U}_{m},\widetilde{V}_{m})$ has the same quality as the solution $(s,t)$ to (9) provided that $(\theta_{i},\tilde{u}_{i},\tilde{v}_{i})$, $i=2,\dots,m$ have comparable accuracy to $(\theta_{1},\tilde{u}_{1},\tilde{v}_{1})$. The resulting method is the basic IPJDSVD, and can also be regarded as a new JDSVD method with the correction equation (56) solved at each outer iteration.
@@
With an approximate solution $(\tilde{s}^{T},\tilde{t}^{T})^{T}$ of (56) available, we orthonormalize $\tilde{s}$ and $\tilde{t}$ against $\widetilde{U}$ and $\widetilde{V}$ to obtain $u_{+}$ and $v_{+}$, and expand the subspaces in a regular way. The above derivation and analysis indicate that the solution to equation (56) expands the searching subspaces as effectively as the solution to the standard correction equation (9) does.
4.2 Convergence analysis on MINRES for solving (56)
@@
(57) $\widetilde{W}_{m}=\begin{bmatrix}\widetilde{P}_{m}&\\ &\widetilde{Q}_{m}\end{bmatrix}\in\mathbb{R}^{(M+N)\times(M+N-2m)}.$
Then the coefficient matrix in (56) can be written as
(58) $\widetilde{B}_{m}=\widetilde{W}_{m}\widetilde{B}_{m}^{\prime}\widetilde{W}_{m}^{T}\quad\mbox{with}\quad\widetilde{B}_{m}^{\prime}=\begin{bmatrix}-\tau I_{M-m}&\widetilde{P}_{m}^{T}A\widetilde{Q}_{m}\\ \widetilde{Q}_{m}^{T}A^{T}\widetilde{P}_{m}&-\tau I_{N-m}\end{bmatrix}.$
Denote by $\tilde{w}_{j}$ the $j$th approximate solution computed by MINRES for (56) with $\tilde{w}_{0}=\bm{0}$, $j=1,2,\ldots,M+N-2m$. Then the corresponding inner residuals are
(59) $\qquad r_{\mathrm{in},j}=\widetilde{B}_{m}\tilde{w}_{j}+r,\qquad\qquad j=0,1,\ldots,M+N-2m.$
@@
With $\widetilde{B}_{m}^{\prime}$ defined in (58), the solution to the preconditioned correction equation (56) is $w=\widetilde{W}_{m}z$ where $z=\widetilde{W}_{m}^{T}w\in\mathbb{R}^{M+N-2m}$ solves
(60) $\widetilde{B}_{m}^{\prime}z=-r^{\prime}\qquad\mbox{with}\qquad r^{\prime}=\widetilde{W}_{m}^{T}r.$
If MINRES is applied to (60) with the initial guess $\tilde{z}_{0}=\bm{0}$, then the $j$th approximate solution $\tilde{z}_{j}=\widetilde{W}_{m}^{T}\tilde{w}_{j}$ with $\tilde{w}_{j}=\widetilde{W}_{m}\tilde{z}_{j}$ and
(61) $\|r_{\mathrm{in},j}\|=\|r^{\prime}_{\mathrm{in},j}\|=\|\widetilde{B}_{m}^{\prime}\tilde{z}_{j}+r^{\prime}\|,\ j=1,2,\ldots,M+N-2m.$
Theorem 4.1 establishes the equivalence of MINRES for solving (56) and (60). Notice that $r=\mathbf{0}$ when $\widetilde{U}_{m}=U_{m}$ and $\widetilde{V}_{m}=V_{m}$. In this case, we can extend Theorem 3.4 to MINRES for solving (56) by supposing any virtual right-hand side $r\not=\mathbf{0}$ that satisfies (51).
@@
Suppose that $\mathcal{R}(\widetilde{U}_{m})=\mathcal{R}(U_{m})$ and $\mathcal{R}(\widetilde{V}_{m})=\mathcal{R}(V_{m})$ in (56) with $U_{m}$ and $V_{m}$ in (48) and $\mathcal{R}(\cdot)$ the column space of a matrix. Then (i) if $\tau>(\sigma_{\max}+\sigma_{\max,m+1})/2$ with $\sigma_{\max,m+1}$ the $(m+1)$th largest singular value of $A$, then
(62) $\frac{\|r_{\mathrm{in},j}\|}{\|r\|}\leq 2\Bigg{(}1-\frac{2}{1+\sqrt{(\tau+\sigma_{\max,m+1})/(\tau-\sigma_{\max,m+1})}}\Bigg{)}^{j};$
(ii) if $\sigma_{\min}\leq\tau<(\sigma_{\min}+\sigma_{\min,m+1})/2$ with $\sigma_{\min,m+1}$ the $(m+1)$th smallest singular value of $A$ and writing $j_{o}=\min\{1,M-N\}$, then
(63) $\frac{\|r_{\mathrm{in},j}\|}{\|r\|}\leq 2\left(\frac{\sigma_{\max}}{\tau}\right)^{j_{o}}\Bigg{(}1-\frac{2}{1+\sqrt{(\sigma_{\max}^{2}-\tau^{2})/(\sigma_{\min,m+1}^{2}-\tau^{2})}}\Bigg{)}^{\left[\frac{j-j_{o}}{2}\right]};$